# Defense Strategies for Autonomous Multi-agent Systems: Ensuring Safety and Resilience Under Exponentially Unbounded FDI Attacks

### v1

@@ Abstract
False data injection (FDI) attacks pose a significant threat to autonomous multi-agent systems (MASs). While resilient control strategies address FDI attacks, they typically have strict assumptions on the attack signals and overlook safety constraints, such as collision avoidance. In practical applications, leader agents equipped with advanced sensors or weaponry span a safe region to guide heterogeneous follower agents, ensuring coordinated operations while addressing collision avoidance to prevent financial losses and mission failures. This letter addresses these gaps by introducing and studying the safety-aware and attack-resilient (SAAR) control problem under exponentially unbounded FDI (EU-FDI) attacks. Specifically, a novel attack-resilient observer layer (OL) is first designed to defend against EU-FDI attacks on the OL. Then, by solving an optimization problem using the quadratic programming (QP), the safety constraints for collision avoidance are further integrated into the SAAR controller design to prevent collisions among followers. An attack-resilient compensational signal is finally designed to mitigate the adverse effects caused by the EU-FDI attack on control input layer (CIL). Rigorous Lyapunov-based stability analysis certifies the SAAR controller’s effectiveness in ensuring both safety and resilience. This study also pioneers a three-dimensional simulation of the SAAR containment control problem for autonomous MASs, demonstrating its applicability in realistic multi-agent scenarios.
Index Terms:
Containment, resilience, unbounded attacks, safety constraints.
@@ I Introduction
Containment control in autonomous multi-agent systems (MASs) has become essential in real-world applications, particularly involving unmanned aerial vehicles (UAVs) [1] and unmanned ground vehicles (UGVs) [2]. This control strategy enables leader agents, often equipped with advanced sensors or weaponry, to guide follower agents within a designated safe region, ensuring coordinated movement and operational safety. Such scenarios are common in tasks such as surveillance, reconnaissance, and military operations, where maintaining cohesion and mitigating risks are critical. The heterogeneity of agents, characterized by differing models, structures, or capabilities, adds complexity to the problem, necessitating advanced control algorithms. Moreover, collision avoidance plays a pivotal role in containment control, as collisions can lead to substantial financial losses and mission failures, highlighting the importance of integrating safety measures into control strategies.
@@
In the MAS framework, false data injection (FDI) attacks pose significant threats to autonomous MASs and impact critical infrastructures and mission-essential cyber-physical systems, such as power systems, transportation networks, and combat-zone multi-robot systems [3]. To mitigate the impact of FDI attacks, two primary approaches have been proposed. The first one involves detection and identification of compromised agents, followed by their removal [4, 5]. However, this method relies on additional assumptions, such as limiting the number of compromised agents. These assumptions are often impractical, attackers typically aim to compromise all agents and communication channels whenever feasible. To address these limitations, a second approach has been developed, focusing on the design of attack-resilient control protocols [6, 7, 8, 9, 10]. Rather than removing compromised agents, this strategy aims to minimize the adverse effects of attacks through resilient control mechanisms. However, the aforementioned research neglects safety constraints, such as collision avoidance, leading to potential collision risks. On the other hand, methods incorporating safety constraints often solve optimization problems for control input but typically lack resilience to FDI attacks [11, 12, 13, 14, 15]. Additionally, most studies addressing FDI resilience assume the observer layer remains intact and either disregard attacks or handle only bounded attack signals with bounded first-time derivatives, which are impractical assumptions [16, 6]. Adversaries can exploit these limitations using rapidly growing signals, such as exponentially unbounded FDI (EU-FDI) attacks, to compromise the system.
@@
In real-world scenarios, ensuring both safety and resilience to EU-FDI attacks is critical for the reliable operation of safety-critical MASs. Recognizing this need, the safe and attack-resilient (SAAR) control problem is studied in this work. The primary challenge lies in designing a controller that is resilient to EU-FDI attacks while consistently enforcing safety constraints, specifically collision avoidance. The contributions are as follows.
@@
$\bullet$ A novel attack-resilient OL is first designed to defend against EU-FDI attacks on the OL. Then, by solving an optimization problem using the QP, the safety constraints for collision avoidance are further integrated into the SAAR controller design to prevent collisions among followers. An attack-resilient compensational signal is finally designed to mitigate the adverse effects caused by the EU-FDI attack on CIL. To the best of the authors’ knowledge, this letter is the first to address EU-FDI attacks on both OL and CIL. Moreover, this letter is the first to investigate both attack-resilience and collision avoidance among followers.
@@
$\bullet$ Rigorous mathematical proof using Lyapunov-based stability analysis is provided to certify the effectiveness of the proposed SAAR controller in guaranteeing both safety constraints and attack resilience under EU-FDI attacks on both OL and CIL.
$\bullet$ This work is the first to simulate a three-dimensional space to investigate the SAAR control strategies, extending the dimensional scope beyond previous two-dimensional space, which reflects a more realistic application for the containment control problem.
This paper is structured as follows: Section II provides the preliminaries and problem formulation. Section III describes the safe and resilient controller design, including the design of the observer, the control strategies, the safety constraints, the main result and we provide the stability analysis and prove that the proposed control guarantees the containment and safety of the autonomous MAS. Finally, Section IV presents the simulation results, and Section V concludes the paper.
@@ II Preliminaries and Problem Formulation
$\|x\|$ denotes the Euclidean norm of a vector $x\in\mathbb{R}^{n}$. The symbol $\partial S$ represents the boundary of a closed set $S$. $L_{f}V(x)=\frac{\partial V}{\partial x}f(x)$ is the Lie derivative of a function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ along a vector field $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. ${I_{N}}\in{\mathbb{R}^{N\times N}}$ is the identity matrix. ${\mathbf{1}}_{N},{\mathbf{0}}_{N}\in{\mathbb{R}^{N}}$ are the column vectors with all elements of zero and one, respectively. The Kronecker product is represented by $\otimes$. The operator $\operatorname{diag}(\cdot)$ forms a block diagonal matrix from its argument. The notations ${\sigma_{\min}}(X)$ and ${\sigma_{\max}}(X)$ represent the minimum singular value and the maximum singular value respectively. $\bar{\square}$ denotes corrupted signal. Consider a system of $N+M$ agents in a digraph $\mathscr{G}$, composed of $N$ followers and $M$ leaders. Let the sets of followers and leaders be denoted by =$\mathscr{F}=\{v_{1},v_{2},\ldots,v_{N}\}$ and =$\mathscr{L}=\{v_{N+1},v_{N+2},\ldots,v_{N+M}\}$, respectively. $\mathcal{N}_{i}$ represents the set of neighboring followers of follower $i$ within the follower set. The interactions between followers are described by the subgraph $\mathscr{G}_{f}=(\mathcal{V},\mathcal{E},\mathcal{A})$, where $\mathcal{V}$ is the set of nodes, $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$ is the set of edges, and $\mathcal{A}=[a_{ij}]\in\mathbb{R}^{N\times N}$ is the adjacency matrix. Here, $a_{ij}$ represents the weight of edge $(v_{j},v_{i})$, with $a_{ij}\neq 0$ if $(v_{j},v_{i})\in\mathcal{E}$, and $a_{ij}=0$ otherwise. There are no repeated edges or self-loops, i.e., $a_{ii}=0$ for all $i$. A directed path from node $i$ to node $j$ is defined as a sequence of edges $\{(v_{i},v_{k}),(v_{k},v_{l}),\ldots,(v_{m},v_{j})\}$. The matrix $\mathcal{G}_{r}=\operatorname{diag}(g_{ir})\in\mathbb{R}^{N\times N}$, where $i\in\mathscr{F}$ and $r\in\mathscr{L}$, represents the diagonal matrix of pinning gains from the $r$-th leader to each follower. Specifically, $g_{ir}\neq 0$ if a connection from the $r$-th leader to the $i$-th follower exists; otherwise, $g_{ir}=0$. The digraph $\mathscr{G}$ is assumed to be time-invariant, meaning that both $\mathcal{A}$ and $\mathcal{G}_{r}$ are constant. $\mathcal{D}=\operatorname{diag}(d_{i})\in\mathbb{R}^{N\times N}$ with $d_{i}=\sum_{j\in\mathcal{N}_{i}}a_{ij}>0$. $\mathcal{L}=\mathcal{D}-\mathcal{A}$. Denote $\Phi_{r}=\frac{1}{M}{\mathcal{L}}+{\mathcal{G}}_{r}$. The states of the leaders are represented by the set $X_{\mathscr{L}}=\{x_{N+1},x_{N+2},\ldots,x_{N+M}\}$.
We consider a group of $N$ heterogeneous agents (followers) described by the following linear dynamics:
$\dot{x}_{i}(t)=A_{i}x_{i}(t)+B_{i}\bar{u}_{i}(t),\quad i\in\mathscr{F},$ (1)
where $x_{i}(t)\in\mathbb{R}^{n}$ is the state, $\bar{u}_{i}(t)\in\mathbb{R}^{m_{i}}$ is the corrupted control input. The matrices $A_{i}\in\mathbb{R}^{n\times n}$ and $B_{i}\in\mathbb{R}^{n\times m_{i}}$ represent the system dynamics and input matrices, respectively. The corrupted input ${\bar{u}_{i}}$ of the $i^{th}$ follower is
$\begin{gathered}\bar{u}_{i}(t)=u_{i}(t)+\gamma^{a}_{i}(t),\hfill\end{gathered}$ (2)
where $u_{i}(t)\in\mathbb{R}^{m_{i}}$ is the intact control input and $\gamma^{a}_{i}(t)\in\mathbb{R}^{m_{i}}$ is of class $C^{1}$ [17] that represents the EU-FDI attack signal injected to the $i^{th}$ follower. The $M$ leaders with the following dynamics can be viewed as command generators that generate the desired trajectories:
@@
where $x_{r}(t)\in\mathbb{R}^{n}$ is the state of the $r$th leader. The matrices $A_{i}$ and $B_{i}$ may vary across different agents, making the system heterogeneous.
@@
$\text{Co}(X_{\mathscr{L}})=\left\{\sum_{r\in\mathscr{L}}a_{r}x_{r}\mid a_{r}\geqslant 0,\sum_{r\in\mathscr{L}}a_{r}=1\right\},$
@@
$S$ has non-repeated eigenvalues on the imaginary axis.
@@
The attack signals $\gamma^{a}_{i}(t)$ and $\gamma^{ol}_{i}(t)$ (in (10)) are exponentially unbounded. That is, their norms grow at most exponentially with time. For the purposes of stability analysis, it is reasonable to assume that there exist positive constants $\kappa^{a}_{i}$ and $\kappa^{ol}_{i}$, such that $\|\gamma^{a}_{i}(t)\|\leqslant\exp(\kappa^{a}_{i}t)$ and $\|\gamma^{ol}_{i}(t)\|\leqslant\exp(\kappa^{ol}_{i}t)$, where $\kappa^{a}$ and $\kappa^{ol}$ could be unknown.
@@
Assumption 4 covers a broad range of FDI attack signals, including those that grow exponentially over time. Specifically, $\exp(\kappa^{a}_{i}t)$ and $\exp(\kappa^{ol}_{i}t)$ represent the worst-case scenarios the controller can handle. As long as the growth rate of the attack signals remains below these thresholds, the controller can effectively mitigate them. In practice, adversaries can inject any time-varying signal into systems through platforms such as software, CPUs, or DSPs. However, much of the existing research focuses on disturbances, noise, or bounded attack signals, or assumes that the first time derivatives of the attacks are bounded [6].
@@
$\begin{gathered}e_{c}=x-\left(\sum_{\nu\in\mathscr{L}}\left(\Phi_{\nu}\otimes I_{n}\right)\right)^{-1}\sum_{r\in\mathscr{L}}\left(\Phi_{r}\otimes I_{n}\right)\bar{x}_{r},\end{gathered}$ (6)
@@
We need the following definition to introduce the concept of a safe set.
@@
The forward invariance of the safe set is guaranteed using control barrier functions (CBFs) . A CBF ensures that the system remains within the safe set by satisfying certain conditions, as described in Definition 4 as follows.
@@
Let $\alpha$ be a monotonically increasing, locally Lipschitz class $\mathcal{K}$ function with $\alpha(0)=0$, given a set $S\subset\mathbb{R}^{n}$ as defined in Definition 3, a function $h(x)$ is CBF candiate, if the following condition is satisfied:
$\inf_{u\in\mathcal{U}}\left[L_{f}h(x)+L_{g}h(x)u\right]\leqslant-\alpha(h(x)),$ (7)
@@
In (7), $L_{f}h(x)$ and $L_{g}h(x)$ denote the Lie derivatives of the function $h(x)$ along the vector fields $f$ and $g$, respectively. For a general system of the form:
$\dot{x}=f(x)+g(x)u,$
the terms $L_{f}h(x)=\frac{\partial h}{\partial x}f(x)$ and $L_{g}h(x)=\frac{\partial h}{\partial x}g(x)$ involve the gradients of $h(x)$ with respect to the state $x$.
@@
$S_{S_{ij}}\triangleq\{x_{i}\in\mathbb{R}^{n}\mid h_{S_{ij}}(x_{i},x_{j})\leqslant 0\},$
@@
$h_{S_{ij}}(x_{i},x_{j})\triangleq d_{s}^{2}-\|x_{i}-x_{j}\|^{2},$
where $d_{s}$ is the minimum allowable safe distance. The system dynamics are given by:
$\dot{x}_{i}(t)=A_{i}x_{i}(t)+B_{i}u_{i}(t),\quad\forall i\in\mathscr{F}.$
In this case, the vector field $f(x)$, applied to linear time invariant system locally, corresponds to $A_{i}x_{i}$, and $g(x)u$ corresponds to $B_{i}u_{i}$. To avoid confusion, we denote the Lie derivatives as $L_{f}h_{S_{ij}}$ and $L_{g}h_{S_{ij}}$, where $f$ and $g$ are defined as above.
Lemma 2.
$\inf_{u_{i}\in\mathcal{U}_{i}}\left[L_{f}h_{S_{ij}}+L_{g}h_{S_{ij}}u_{i}\right]\leqslant-\alpha(h_{S_{ij}})-2(x_{i}-x_{j})^{\top}\dot{x}_{j},$ (8)
Proof: By definition, the specific CBF is given as:
$h_{S_{ij}}(x_{i},x_{j})=d_{s}^{2}-\|x_{i}-x_{j}\|^{2}.$
Taking the time derivative of $h_{S_{ij}}(x_{i},x_{j})$ along the system dynamics, we have:
$\dot{h}_{S_{ij}}(x_{i},x_{j})=-\frac{d}{dt}\|x_{i}-x_{j}\|^{2}.$
Expanding $\|x_{i}-x_{j}\|^{2}=(x_{i}-x_{j})^{\top}(x_{i}-x_{j})$, its time derivative is:
$\frac{d}{dt}\|x_{i}-x_{j}\|^{2}=2(x_{i}-x_{j})^{\top}(\dot{x}_{i}-\dot{x}_{j}).$
Substituting this into $\dot{h}_{S_{ij}}(x_{i},x_{j})$, we obtain:
$\dot{h}_{S_{ij}}(x_{i},x_{j})=-2(x_{i}-x_{j})^{\top}(\dot{x}_{i}-\dot{x}_{j}).$
From the system dynamics, we know:
$\dot{x}_{i}=A_{i}x_{i}+B_{i}u_{i},\quad\dot{x}_{j}=A_{j}x_{j}+B_{j}u_{j}.$
Substituting $\dot{x}_{i}$ and $\dot{x}_{j}$ into the expression for $\dot{h}_{S_{ij}}(x_{i},x_{j})$, we get: $\dot{h}_{S_{ij}}(x_{i},x_{j})=-2(x_{i}-x_{j})^{\top}\left[(A_{i}x_{i}+B_{i}u_{i})-(A_{j}x_{j}+B_{j}u_{j})\right].$
Rearranging terms, we separate the contributions of $x_{i}$, $x_{j}$, and $u_{i}$:
$\begin{gathered}\dot{h}_{S_{ij}}(x_{i},x_{j})=-2(x_{i}-x_{j})^{\top}A_{i}x_{i}+2(x_{i}-x_{j})^{\top}A_{j}x_{j}\\ -2(x_{i}-x_{j})^{\top}B_{i}u_{i}+2(x_{i}-x_{j})^{\top}B_{j}u_{j}.\end{gathered}$ (9)
Using the Lie derivative notation, this can be written as:
$\dot{h}_{S_{ij}}(x_{i},x_{j})=L_{f}h_{S_{ij}}+L_{g}h_{S_{ij}}u_{i}-2(x_{i}-x_{j})^{\top}\dot{x}_{j},$
where, $L_{f}h_{S_{ij}}=-2(x_{i}-x_{j})^{\top}A_{i}x_{i},L_{g}h_{S_{ij}}=-2(x_{i}-x_{j})^{\top}B_{i}.$
To ensure forward invariance of $S_{S_{ij}}$, we impose the condition:
$\dot{h}_{S_{S_{ij}}}(x_{i},x_{j})\leqslant-\alpha(h_{S_{ij}}).$
Substituting the expanded form of $\dot{h}_{S_{ij}}(x_{i},x_{j})$, we obtain:
$L_{f}h_{S_{ij}}+L_{g}h_{S_{ij}}u_{i}\leqslant-\alpha(h_{S_{ij}})-2(x_{i}-x_{j})^{\top}\dot{x}_{j}.$
This inequality ensures that the safe set $S_{S_{ij}}$ is forward invariant. Thus, if the control input $u_{i}$ satisfies the optimization criterion:
$\inf_{u_{i}\in\mathcal{U}_{i}}\left[L_{f}h_{S_{ij}}+L_{g}h_{S_{ij}}u_{i}\right]\leqslant-\alpha(h_{S_{ij}})-2(x_{i}-x_{j})^{\top}\dot{x}_{j},$
then $\|x_{i}-x_{j}\|\geqslant d_{s}$ is guaranteed for all $t\geqslant 0.\quad\quad\quad\quad\blacksquare$
@@
Definition 5 (SAAR).
For the heterogeneous autonomous MAS described in (1)-(3) under EU-FDI attacks, the SAAR control problem is to design control input $u_{i}\in\mathcal{U}_{i}$ in (2), such that 1) the UUB containment control objective is achieved, i.e., the state of each follower converges to a small neighborhood around or within the dynamic convex hull spanned by the states of the leaders and 2) the state $x_{i}(t)$ remains within the safe set $S_{S_{ij}}$ for all $t\geqslant 0$ in the presence of EU-FDI attacks on both OL and CIL.
Remark 2.
The SAAR problem addresses key challenges in practical applications of autonomous MASs. In scenarios such as surveillance, reconnaissance, and military operations, leader agents (e.g., UAVs or UGVs) equipped with advanced sensors or weaponry span a safe region to guide follower agents, ensuring coordinated and secure operations. The heterogeneity of these systems, characterized by diverse agent models and dynamics, necessitates control strategies that accommodate such differences. Moreover, collision avoidance is integral to maintaining operational safety and preventing financial and mission-critical losses. These considerations naturally align with the requirements addressed in the SAAR problem, demonstrating its relevance to real-world MASs.
III Safe and Resilient Controller Design
In this section, we propose a fully distributed, safe, and resilient containment control framework to address the SAAR control problem. First, we introduce a fully distributed, attack-resilient OL to estimate the convex combinations of the leaders’ states.

### v2

# Certified Resilient Collision-free Control Framework for Heterogeneous Multi-agent Systems Against Exponentially Unbounded FDI Attacks

@@ Abstract
False data injection attacks pose a significant threat to autonomous multi-agent systems (MASs). Existing attack-resilient control strategies generally have strict assumptions on the attack signals and overlook safety constraints, such as collision avoidance. In practical applications, leader agents equipped with advanced sensors or weaponry span a safe region to guide heterogeneous follower agents, ensuring coordinated operations while addressing collision avoidance to prevent financial losses and mission failures. This letter addresses these gaps by introducing and solving the safety-aware and attack-resilient (SAAR) control problem under exponentially unbounded false data injection (EU-FDI) attacks. Specifically, a novel attack-resilient observer layer (OL) is first designed to defend against EU-FDI attacks on the OL. Then, an attack-resilient compensational signal is designed to mitigate the adverse effects caused by the EU-FDI attack on control input layer (CIL). Finally, a SAAR controller is designed by solving a quadratic programming (QP) problem integrating control barrier function (CBF) certified collision-free safety constraints. Rigorous Lyapunov-based stability analysis certifies the SAAR controller’s effectiveness in ensuring both safety and resilience. This study also pioneers a three-dimensional (3D) simulation of the SAAR containment control problem for heterogeneous MASs, demonstrating its applicability in realistic multi-agent scenarios.
@@ I Introduction
Containment control in autonomous MASs has become essential in real-world applications, particularly involving unmanned aerial vehicles (UAVs) [1] and unmanned ground vehicles (UGVs) [2]. This control strategy enables leader agents, often equipped with advanced sensors or weaponry, to guide follower agents within a designated safe region, ensuring coordinated movement and operational safety. Such scenarios are common in tasks such as surveillance, reconnaissance, and military operations, where maintaining cohesion and mitigating risks are critical. The heterogeneity of agents, characterized by differing models, structures, or capabilities, adds complexity to the problem, necessitating advanced control algorithms. Moreover, collision avoidance plays a pivotal role in containment control, as collisions can lead to substantial financial losses and mission failures, highlighting the importance of integrating safety measures into control strategies.
@@
In the MAS framework, FDI attacks pose significant threats to autonomous MASs and impact critical infrastructures and mission-essential cyber-physical systems, such as power systems, transportation networks, and combat-zone multi-robot systems [3]. To mitigate the impact of FDI attacks, two primary approaches have been proposed. The first one involves detection and identification of compromised agents, followed by the attack-isolation [4, 5]. However, this method relies on strict limitations, such as the number of compromised agents, which are often impractical. To address these limitations, a second approach has been developed, focusing on the design of attack-resilient control protocols [6, 7, 8, 9, 10]. Rather than detecting and removing compromised agents, this strategy aims to minimize the adverse effects of attacks through resilient control mechanisms. However, the aforementioned research neglects safety constraints, such as collision avoidance, leading to potential collision risks. On the other hand, existing safe control methods incorporating safety constraints often fail to guarantee resilience for MASs against FDI attacks [11, 12, 13, 14, 15]. Additionally, most studies addressing FDI resilience for heterogeneous MASs assume the OL remains intact and either disregard attacks or handle only bounded attack signals or the ones with bounded first-time derivatives, which are impractical assumptions [16, 7]. However, malicious adversaries could inject any time-varying signals, such as EU-FDI attacks, to compromise the system.
@@
In real-world scenarios, ensuring both safety and resilience against EU-FDI attacks is crucial for the reliable operation of safety-critical MASs. To address this challenge, this letter formulates and solves the SAAR control problem. The key difficulty lies in designing a controller that effectively mitigates EU-FDI attacks while consistently enforcing collision avoidance constraints. The main contributions are as follows.
@@
$\bullet$ A novel attack-resilient OL is first designed to defend against EU-FDI attacks on OL. Then, an attack-resilient compensation signal is developed to mitigate the adverse effects of EU-FDI attacks on CIL. Finally, the SAAR controller is designed by solving a QP problem incorporating CBF-certified collision-free safety constraints. To the best of the authors’ knowledge, this letter is the first to guarantee both resilience and safety for heterogeneous MASs, in the presence of EU-FDI attacks on both OL and CIL.
@@
$\bullet$ A rigorous Lyapunov-based stability analysis is conducted to certify the effectiveness of the proposed SAAR controller in ensuring both safety and resilience. Its efficacy is further demonstrated through a 3D simulation of the containment control problem for heterogeneous MASs.
@@ II Preliminaries and Problem Formulation
$\|x\|$ denotes the Euclidean norm of a vector $x\in\mathbb{R}^{n}$.
$L_{f}V(x)=\frac{\partial V}{\partial x}f(x)$ is the Lie derivative of a function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}$ along a vector field $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. ${I_{N}}\in{\mathbb{R}^{N\times N}}$ is the identity matrix. ${\mathbf{1}}_{N},{\mathbf{0}}_{N}\in{\mathbb{R}^{N}}$ are the column vectors with all elements of zero and one, respectively. The Kronecker product is represented by $\otimes$. The operator $\operatorname{diag}(\cdot)$ forms a block diagonal matrix from its argument. The notations ${\sigma_{\min}}(\cdot)$ and ${\sigma_{\max}}(\cdot)$ represent the minimum singular value and the maximum singular value respectively. Consider a system of $N+M$ agents on a time-invariant digraph $\mathscr{G}$, consisting of $N$ followers and $M$ leaders, denoted by =$\mathscr{F}=\{v_{1},v_{2},\ldots,v_{N}\}$ and =$\mathscr{L}=\{v_{N+1},v_{N+2},\ldots,v_{N+M}\}$, respectively. $\mathcal{N}_{i}$ represents the set of neighboring followers of follower $i$. The interactions among followers are described by the subgraph $\mathscr{G}_{f}=(\mathcal{V},\mathcal{E},\mathcal{A})$. More details on the digraph can be found in [6]. Denote $\Phi_{r}=\frac{1}{M}{\mathcal{L}}+{\mathcal{G}}_{r}$. The states of the leaders are represented by the set $X_{\mathscr{L}}=\{x_{N+1},x_{N+2},\ldots,x_{N+M}\}$.
We consider a group of $N$ heterogeneous followers described by the following dynamics:
$\dot{x}_{i}(t)=A_{i}x_{i}(t)+B_{i}u_{i}(t),\quad i\in\mathscr{F},$ (1)
where $x_{i}(t)\in\mathbb{R}^{n}$ is the state, $u_{i}(t)\in\mathbb{R}^{m_{i}}$ is the control input. The system matrices $A_{i}$ and $B_{i}$ may vary across different agents, making the system heterogeneous. The $M$ leaders with the following dynamics can be viewed as command generators that generate the desired trajectories:
@@
where $x_{r}(t)\in\mathbb{R}^{n}$ is the state of the $r$th leader.
@@
$\text{Co}(X_{\mathscr{L}})=\left\{\sum\nolimits_{r\in\mathscr{L}}a_{r}x_{r}\;\middle|\;a_{r}\geqslant 0,\sum\nolimits_{r\in\mathscr{L}}a_{r}=1\right\}.$
@@
$S$ has eigenvalues with non-positive real parts, and non-repeated eigenvalues on the imaginary axis.
@@
The FDI attack signals on the CIL and OL, $\gamma^{a}_{i}(t)$ and $\gamma^{ol}_{i}(t)$, are exponentially unbounded. That is, their norms grow at most exponentially with time. For the purposes of stability analysis, it is reasonable to assume that there exist positive constants $\kappa^{a}_{i}$ and $\kappa^{ol}_{i}$, such that $\|\gamma^{a}_{i}(t)\|\leqslant\exp(\kappa^{a}_{i}t)$ and $\|\gamma^{ol}_{i}(t)\|\leqslant\exp(\kappa^{ol}_{i}t)$, where $\kappa^{a}_{i}$ and $\kappa^{ol}_{i}$ could be unknown.
@@
Assumption 4 covers a broad range of FDI attack signals, including those that grow exponentially over time. Specifically, $\exp(\kappa^{a}_{i}t)$ and $\exp(\kappa^{ol}_{i}t)$ represent the worst-case scenarios the controller can handle. As long as the attack signals remain below these envelops, the controller can effectively mitigate them. In practice, adversaries can inject any time-varying signal into MAS through platforms such as software, CPUs, or DSPs. However, most of the existing research focuses on disturbances, noise, or bounded attack signals, or assumes that the first time derivatives of the attacks are bounded [7].
@@
$\begin{gathered}e_{c}=x-\left(\sum\nolimits_{\nu\in\mathscr{L}}\left(\Phi_{\nu}\otimes I_{n}\right)\right)^{-1}\sum\nolimits_{r\in\mathscr{L}}\left(\Phi_{r}\otimes I_{n}\right)\bar{x}_{r},\end{gathered}$ (5)
@@
The forward invariance of the safe set is guaranteed using CBFs, as described as follows.
@@
Let $\alpha$ be a monotonically increasing, locally Lipschitz class $\mathcal{K}$ function with $\alpha(0)=0$, given a set $S\subset\mathbb{R}^{n}$ as defined in Definition 3, a function $h(x)$ is CBF, if the following condition is satisfied:
$\inf_{u\in\mathcal{U}}\left[L_{f_{i}}h(x)+L_{g_{i}}h(x)u\right]\leqslant-\alpha(h(x)),$ (6)
@@
In (6), $L_{f}h(x)$ and $L_{g}h(x)$ denote the Lie derivatives of the function $h(x)$ along the vector fields $f$ and $g$, respectively. For a general system of the form: $\dot{x}=f(x)+g(x)u,$ the terms $L_{f}h(x)=\frac{\partial h}{\partial x}f(x)$ and $L_{g}h(x)=\frac{\partial h}{\partial x}g(x)$ involve the gradients of $h(x)$ with respect to the state $x$. if control input $u$ is designed such that (6) is satisfied, the forward invariant set $S$ is obtained [21].
@@
$S_{S_{ij}}\triangleq\{x_{i}\in\mathbb{R}^{n}\mid h_{S_{ij}}(x_{i},x_{j})\leqslant 0\},$ (7)
@@
$h_{S_{ij}}(x_{i},x_{j})\triangleq d_{s}^{2}-\|x_{i}-x_{j}\|^{2},$ (8)
where $d_{s}$ is the minimum allowable safe distance. The system dynamics are given by: $\dot{x}_{i}(t)=A_{i}x_{i}(t)+B_{i}u_{i}(t),\,\forall i\in\mathscr{F}.$ In this case, the vector field $f_{i}(x)$, applied to linear time invariant system locally, corresponds to $A_{i}x_{i}$, and $g_{i}(x)$ corresponds to $B_{i}$. To avoid confusion, we denote the Lie derivatives as $L_{f_{i}}h_{S_{ij}}$ and $L_{g_{i}}h_{S_{ij}}$, where $f_{i}$ and $g_{i}$ are defined as above.
@@
Definition 5 (SAAR Control Problem).
For the heterogeneous MAS described in (1)-(2) under EU-FDI attacks, the SAAR control problem is to design control input $u_{i}\in\mathcal{U}_{i}$, such that 1) $e_{c}$ is UUB, i.e., the UUB containment control objective is achieved. Specifically, the state of each follower converges to a small neighborhood around or within the dynamic convex hull spanned by the states of the leaders; 2) the state $x_{i}(t)$ remains within the safe set $S_{S_{ij}}$ for all $t\geqslant 0$ in the presence of EU-FDI attacks on both OL and CIL.
III Main Result of SAAR Controller Design
In this section, we propose a fully distributed, safe, and resilient containment control framework to address the SAAR control problem. Since only the neighboring followers of the leaders have access to the leaders’ states, to achieve the containment control objective for each follower, a OL dynamics is needed to estimate the convex combinations of the leaders’ states. Here, we introduce the following fully distributed and attack-resilient OL dynamics